OpenStudy (anonymous):

how to evaluate negatine 7 over 10 divided by 11 over 10

OpenStudy (anonymous):

\[-\frac 7{10} \div \frac{11}{10} = -\frac{7}{10} \times \frac{10}{11}\]

OpenStudy (anonymous):

Can you do it now?

OpenStudy (anonymous):

no they dont have a common denominator

OpenStudy (anonymous):

You're multiplying. Not adding or subtracting.

OpenStudy (anonymous):

A common denominator is unnecessary.